You are serving on a jury. A plaintiff is suing the city for injuries sustained after a freak street sweeper accident. In the trial, doctors testified that it will be five years before the plaintiff is able to return to work. The jury has already decided in favor of the plaintiff. You are the foreperson of the jury and propose that the jury give the plaintiff an award to cover the following:

(a)

The present value of two years’ back pay. The plaintiff’s annual salary for the last two years would have been $59,000 and $61,000, respectively.

(b)

The present value of five years’ future salary. You assume the salary will be $64,000 per year.

(c) $200,000 for pain and suffering.
(d)

$20,000 for court costs.

Assume the salary payments are equal amounts paid at the end of each month.
Requirement 1:

If the interest rate you choose is an EAR of 11 percent, what is the size of the settlement? (Enter youranswer as directed, but do not use rounded numbers in intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)
Answer #1

Requirement 1:

a.Present value of 2 years back pay = 59000*1.11^2 + 61000*1.11 = 140,403.90

b. Present value of 5 years future pay =PV(rate,nper,pmt) in excel =PV(0.11,5,64000) = 236,537.41

c. 200,000

d. 20,000

Size of settlement = 140,403.90 + 236,537.41 + 200,000 + 20,000 = $596,941.31